OpenStudy (anonymous):

How do you get the vector equation for the tangent line to the graph of r at t=1 given: r(1)= <-3,0,2> r'(1)=<5,2,0> ||r'(1)||=6 r'''(1)=<0,-8,3>?

OpenStudy (anonymous):

is it ||r''(1)|| = 6 or ||r'(1)||? r' is <5,2,0>

OpenStudy (anonymous):

It's ||r'(1)|| and r'(1) is <5,2,0>

OpenStudy (amistre64):

r' is the direction vector of the line when t=1 r is the point that it is attached to when t=1

OpenStudy (amistre64):

L = r + tr'

OpenStudy (amistre64):

L = r(1) + tr'(1) that is :)

OpenStudy (anonymous):

Thanks. But how would the vector equation be written? Like this: <-3+5t, 2t, 2>?

OpenStudy (amistre64):

that doesnt seem like a vector equation to me. i believe the notation for that form is: given a point P and a vector V, the line L is formed as: L = P + tV

OpenStudy (amistre64):

it can be parametricised as: P=(px,py,pz), V=(vx,vy,vz) L=: x = px + t vx y = py + t vy z = pz + t vz

OpenStudy (anonymous):

So what would be the final answer to the problem?

OpenStudy (amistre64):

L = r(1) + t r'(1) ... as i stated before.

OpenStudy (anonymous):

Oh right! Thanks a bunch! :)

OpenStudy (amistre64):

good luck ;)